1. POP110 RF amps are powered from the cross connect. But that +15V block has GND connections that are not connected to the ground.
i.e. The ground potential is given by the signal ground. (Attachment 1)
This is caused by the misuse of the DIN connector blocks. The hod side uses an isolated block assuming a fuse is inserted.
However, the ground sides also have the isolated blocks
2. One of the POP110 RF cable has a suspicious shiled. The rigidity of the cable is low, suggesting the broken shield. (Attachment 2)
The variable delay line has been setup for practical use. The hardware and basic software are ready.
The delay time is given by [512-1-mod(C1:LSC-BO_1_0_SET, 512)]*(1/16) ns
Giving 511 (LLLL LLLH HHHH HHHH) to C1:LSC-BO_1_0_SET makes the delayline shortest (+0ns).
Giving 0 (LLLL LLLL LLLL LLLL) to C1:LSC-BO_1_0_SET makes the delayline longest (~32ns).
The SR785 was removed from the rack for our access >> Eric
- Three CONTEC DO-32L-PE cards are found in the Yarm digital cabinet. (I brought a card from WB, but will bring it back).
- The card was installed in the C1LSC chassis.
- The models for c1x04 and c1lsc were modified to include the card. Once they are restarted, the card was recognized without problem.
The frame builder also needed to be restarted (Attachment 1&2). The changes were committed to the repository.
- MEDM screen "CDS_BO_STATUS.adl" has been modified to include the bit monitors for the new DO card. (Attachment 3)
Epics values "C1:LSC-BO_1_0_SET" and "C1:LSC-BO_1_1_SET" are hooked up to the DO block.
- The DO board has DB37(F). I made a I/F cable with a DB37(M) crimp connector, DB25 breakout board, and a ribbon cable.
Pin 1 is connected to pin 14 of the DB25 (GND of the delayline circuit).
Pin 2~10 are connected to pin 1~9 of the DB25 (Switch 1~9 of the delayline circuit)
Pin 18 is connected to X01 (external = spare) (Attachment 4)
- [CONFESSION] A bench +15V power supply was prepared to power the transisters of the DO (Attachment 6). The hot side is connected to X01 (not connected to the DB25),
and the cold side is connected to pin 14 of the DB25. Once we find this is a useful setup we need to make a dedicated interface unit to convert DB37
into DB25 (and provide more connectivities).
- A DB25 M-F cable was installed on the cable tray above the LSC racks.
Delay line unit
- The delay line box was mounted on 34H of the LSC analog rack (Attachment 5).
- The side cross connect power supply was not available (to be described later). Therefore we decided to use the same +15V supply as the one for the DO card.
- Checked the functionarity of the local switches using a function generator @30MHz and the front panel switches. The maximum (~32ns) delay was confirmed.
(Just not enough to have 360 deg shift).
- Now the delay line function was tested with the front panel swicth at "ext". We confirmed that the delay time changes with the number given to C1:LSC-BO_1_0_SET.
What we need further
- Implement delay time slider control (511 = 0ns, 0 = 31.94ns). The delay time is given by
[512-1-mod(C1:LSC-BO_1_0_SET, 512)]*(1/16) ns
- Some independent RF issues I found. (Next entry)
About the analog CARM control with ALS:
We're looking at using a Sigg designed remotely switchable delay line box on the currently undelayed side of the ALS DFD beat. For a beat frequency of 50MHz, one cycle is 20ns, this thing has 24ns total delay capability, so we should be able to get pretty close to a zero crossing of the analog I or Q outputs of the demod board. This can be used as IN2 for the common mode board.
Gautam is testing the functionality of the delay and switching, and should post a link to the DCC page of the schematic. Rana and Koji have been discussing the implementation of the remote switching (RCG vs. VME).
I spent some time this afternoon trying to lock the X arm in this way, but instead of at IR resonance, just wherever the I output of the DFD had a zero crossing. However, I didn't give enough thought to the loop shapes; Koji helped me think it through. Tomorrow, I'll make a little pomona box to go before the CM IN2 that will give the ALS loop shape a pole where we expect the CARM coupled cavity pole to be (~120Hz), so that the REFL11 and ALS signals have a similar shape when we're trying to transition.
The common mode board does have a filter for this kind of thing for single arm tests, but puts in a zero as well, as it expects the single arm pole, which isn't present in the ALS sensing, so maybe I'll whip up something appropriate for this, too.
ETMY optical table enclosure feedthrough- south is in. Now it is time to see how air tightness increases performance.
The ETMY enclosure feedthrough - north is installed. The sealing material is hard to work with.
The upper empty blocks will be replaced by something soft to make changing cables easy.
While investigating the BIO situation with the LSC machine and the iscaux2 processor last night, we wondered if maybe the Anti-Aliasing filters were mistakenly disabled. But why do we need these anyway?
Our ADCs digitize at 64 kHz and there is a digital lowpass in the IOP at 5 kHz before we downsample to 16 kHz. So mainly we're trying to prevent some aliasing at the 64 kHz IOP rate. But our analog AA filter is a 8th order ELP at 7570 Hz, so its overkill.
So, I propose that we bypas the AA via hardwiring the board and implement a 10 kHz pole in the whitening board (D990694) before the whitening by turning R127, etc. into a 0.1 uF cap. Along with the 100 Ohm series resistor, this will make a pole at ~15 kHz. Probably ought to check that the input resistor is metal film. Also, if we replace C158/C159, etc. with a 0.47 nF cap, we'll get 2 poles at 35 kHz to limit the higher frequencies from saturating.
Tonight we noticed that the REFL_DC signal has gone bipolar, even though the whitening gain is 0 dB and the whitening filter is requested to be OFF.
We should check out the switch operation of several ofthe LSC channels in the daytime - where is the procedure for this diagnostic posted?
With the adjustable delay line box installed in the 55MHz modulation path, I've measured the PRMI sensing matrix as a function of delay / relative phase between the 11MHz and 55MHz modulations. The relative frequency difference of 44MHz tells us that this should be cyclical after ~23nsec of delay, but losses in the delay cable change this; see Koji's elogs about the modulation cancellation setup for details.
TL;DR: Nothing really changes, other than REFL33 optical gain. MICH/PRCL angles remain degenerate.
The results aren't so surprising. The demod angles for the 55MHz diodes don't even change, since the same 55MHz signal is used for the modulator and demodulators, so delaying it before the split should go unnoticed. Most of these measurements were made during the same lock stretch, PRCL on REFL11 I and MICH on AS55Q.
The only signals we would expect to change much are ones that have significant contriubtions from field products influenced by both modulations. None of the 1F PDs are like this, nor is REFL165. REFL33 is the odd man out, where the +44MHz field produced as a -11MHz sideband on the +55MHz sideband beats with the +11MHz sideband (and the same with the signs flipped). I made a simulation for the 40m poster at the March 2015 LVC meeting, but I don't think it ever made it to the ELOG.
Here are the results for the 0ns and 4ns cases, as an illustration of what changes (REFL33), and what doesn't (everything else). Again, these are calibrated to Volts out of the analog demod boards per meter of DoF motion.
So, since REFL33 is the only one really changing, let's just look at it by itself:
Qualitatively, the change in magnitude looks similar to the simulation result. The demod angles fall by some roughly linear amount. The angle difference is even more stationary than predicted there, though.
I turned on the MCF FF in the OAF today (we need to fix the labeling of the 'ON' buttons on the RHS of the screen). The performance is still good; before / after attached.
Not only is the 1 Hz performance in the MC still good, but the X & Y arm noise reduction is ~1 order of magnitude. Good to know that the filters aren't changing much with time.
Can we just leave this on all the time now? Seems to be OK and there's no visible increase in the arm noise with this on.
Also did some updates to the summary pages and added a CDS FEC tab for CPU times.
Please take a look at the summary pages and bring a list of demands to the Wednesday meeting.
Just a heads up while I'm out for a bit: the delay line is currently installed in the 55MHz modulation path.
I'll be back later, and will revert the setup.
TP2 dry fore pump sn PLE10082 was replaced at pressure 717 mTorr, TP2 50K rpm 0.33A @ 112,677 hrs
Top seal life was 8,160 hrs
Model SH110, Sn LP1007L556 was installed. It's fore line pressure after 30 minutes of running 38 mTorr, TP2 turbo at 50K rpm 0.18A
When making the Wiener filter OFF/ON comparisons, we want to use the median PSD estimates, not the mean (which is what pwelch gives you).
cf. Sujan's note and Evan's follow-up
The median will be less sensitive to the transients / gltiches and will show more improvement I think.
The following MCL filters were left loaded in the T240-X and T240-Y FF filter modules (filters go in pairs, both on):
FM7: SISO filters for MCL elog:11541
FM8: MISO v1 elog:11547
FM9: MISO v1.1 Small improvement over MISO v1
FM10 MISO v2 elog:11563
FM5 MISO v3.1 elog:11584 (best one)
FM6 MISO 3.1.1 elog:11584 (second best one)
POP110 and POP22 demod angles were adjusted for DRMI lock.
Last week, I never achieved a fully 1F lock, REFL165 was used for SRCL. Tonight, we created input matrix settings for pure 1F locking, and did some signal mixing to reduce the PRCL to SRCL coupling. The PRCL to MICH coupling was already low, since AS55 is fairly insensitive to PRCL.
Similarly, for the 3F signals, some signal mixing of REFL33I and REFL165Q was used to reduce the PRCL to MICH coupling. The PRCL to SRCL coupling in REFL165 isn't too bad, so no compensation was done. Interestingly, in this setting, the 3F MICH and SRCL signals agree with the 1F signals on their zero crossing, so no offsets are needed. REFL33 I does need an offset, however, to match the REFL11I PRCL zero crossing.
The DRMI acquires faster with SRCL set to 165I. Once acquired, the 1F/3F can be made smoothly, and both settings are very stable. The sensing matrix in each setting is consistent with each other. (The PRCL and SRCL lines in AS55 change, but really I shouldn't even plot them, since they're not very coherent).
For some reason, these show a sign flip relative to last week's measurements. The relative angles are consistent, though.
Next up is finding the right coefficient for the SRM in the MICH output matrix, when actuating on the BS.
We looked at the DRMI noise spectrum and chose new excitation frequencies such that the lines are lower in frequency than before (~300 Hz instead of 800 Hz) and also not in some noisy region.
New filters is saved and loaded for all LSC DOFs.
Our online subtraction filters for PRC angle and MC length were trained on the raw ADC signals, so I've inverted the filters that Rana installed in the PEM filter banks in the OAF signal conditioning filter banks (C1:OAF-WIT_STS1X, etc.)
It's not perfect, since the inversion would be unstable, and thus needs a low pass. I used an ELP at 800Hz. The error in the inversion is then something like half a degree at 5Hz, which is the highest frequency we really ever subtract at. It should be ok.
On Thursday night (sorry for the late elog) I decided to give the MCL FF one more try.
I first remeasured the actuator transfer function because previous measurements had poor coherence ~0.5 - 0.7 at 3 Hz. I did a sine swept to measure the TF.
Raw transfer function:
The data is attached here: TF.zip
Then I made Wiener filters by fitting the transfer function data with coherence > 0.95 (on the left). Fitting all the data (on the right). Here are the filters:
The offline subtractions (high coh fit on left, all data fit on right). Notice the better IIR performance when all the TF data was fitted.
The online results: (these were aquired by taking five DTT measurements with 15 averages each and then taking the mean of these measurements)
And the subtraction performance:
I converted our MC WFS relief from CSH to BASH today. I also added 'wait' commands and 'echo' commands so that all DoFs run in parallel nicely. It can be accessed from the MC WFS screen.
I increased the overall MC WFS gain input slider from 0.02 to 0.1 (its in the mcwfson script). The MC Trans loops now have a time constant of ~30 seconds. The relief script relieves ~90% of the MC WFS control signals in the 2 minutes that its allowed to run.
On the next upgrade, we should make it python and have it kill the relief process if the MC loses lock before relief is applied via the alignment sliders.
As it turned out, the "STS" BLRMS filters were all a mess, so I fixed them up today:
The "C1:PEM-SEIS_STS_1" filter banks are currently empty, so the signal is just in ADC counts. However, by amazing luck, this seems to be the right gain (within a factor of 2) to put the signal into units of microns / second. According the the schematic (D1000749), the default gain of 110 can be switched to make the whole box just have a gain of 2 (differential in, differential out). I wonder if anyone, like Jenne, knows if this is what we have? There's no elog I found about setting the gain switch.
According to the manual, the gain is ~1175 V/(m/s). Our ADC gain should be (2^16)/(40). So:
cal_gain = 1175 * 2 * 65536 / 40 ==>> 0.26 (m/s)/counts
I have put this into the STS_1_X,Y,Z filter modules in c1pem so that these channels are now calibrated. I also put the first few s-domain poles/zeros into the filter based on the manual so that the magnitude in the 10-30 Hz band is correct-ish now.
* Does anyone know how to center the masses on this thing?
The new stage missed the right height by ~2 mm.
Even if I completely bottom out the (New Focus 9071) 4-axis stage, its not short enough. So I removed the AOM from the beam and re-aligned into the PMC.
Steve, please get the aluminum piece remachined to go down by 2.5 mm so we can have some height adjustment room.
New stage can hold the correct polarization.
Also, the turning mirror mount just after the EOM and before the AOM is a U-100 and we want it to be a Suprema for stability - let's not forget to swap that after Steve gets the mount fixed.
Frustrated by the single pixel width of the windows and how hard that makes it to drag things around, I explored StackExchange:
which showed how there is a .xml file which can be edited to increase this. I've changed the border size to 4 pixels on Rossa - its nice.
Some years ago I bought some dividers from Wenzel. For each arm, we have x256 and a x64 divider. Wired in series, that means we can divide each IR beat by 2^14.
The highest frequency we can read in our digital system is ~8100 Hz. This corresponds to an RF frequency of ~132 MHz which as much as the BBPD could go, but less than the fiber PDs.
Today we checked them out:
Since this seems promising, we're going to make a box on Monday to package both of these. There will one SMA input and output per channel.
Each channel will have a an amplifier since this need not be a low noise channel. The ZKL-1R5 seems like a good choice to me. G=40 dB and +15 dBm output.
Then Gautam will make a frequency counter module in the RCG which can do counting with square waves and not care about the wiggles in the waveform.
I think this ought to do the trick for our Coarse frequency discriminator. Then our Delay Box ought to be able to have a few MHz range and do all of the Fast ALS Carm that we need.
Nice going. I think the LLO / LHO scheme is to acquire on 1F and then cdsutils avg to get the 3F offsets. The thinking is that that 1F signals have less intrinsic offset than the 3F signals, so we want to be use digital offsets for the 3F locks.
I've now made a collection of sensing matrix measurements.
In all of the plots below, the radial scale is logarithmic, each grid line is a factor of 10. The units of the radial direction are calibrated into demod board output Volts per meter. The same radial scale is used on all plots and subplots.
I did two PRMI measurements: with MICH locked and excited with either the ITMS or the BS + PRM compensation. This tells us if our PRM compensation is working; I think it is indeed ok. I though I remembered that we came up with a number for the SRM compensation, but I haven't been able to find it yet.
The CARM sensing int he PRFPMI measurement has the loop gain at the excitation frequency undone. All excitations were simultaneously notched out of all control filters, via the NotchSensMat filters.
The angular scale is set to the analog I and Q signals; the dotted lines show the digitial phase rotation angle used at the time of measurement.
New stage can hold the correct polarization.
DRAWING CORRECTION: Post block height was lowered to be 1.88" from 2.0"
I was thinking that the "FOSC" product line (which is called a "coupler" instead of a "splitter/combiner") was what we wanted.
Koji brought to my attention that the 90/10 splitters we already have are of this line. So, I rigged a few up to shine a hopefully beating pair of fields on the fiber coupled thorlabs PD.
I was able to get ~80uW each of PSL and AUX X light on the PD, which produced a -10dBm beatnote! Thus, I think these FOSC splitters are indeed what we want.
I then threw this IR beatnote at our ALS signal chain. The beatnote was too big to throw through our ~+27dB RF amps, so I just sent the -10dBm over to the LSC rack.
The IR beat spectrum is somwhat noisier from 10-100Hz, but, more interesting, is that the sub-4Hz noise is identical in the two beats, and very coherent. This excludes ALS noise arising from anything happening in the green beat optics on the PSL table.
Obviously, the high frequency noise is largely the same and coherent too, but also coherent with the AUX X PDH control signal, so it is understood.
Single mode coupler, 2x2, 1064nm +/-20nm, 50/50 ratio, 900micron loose tube jacket, Hi1060flex fiber, 1m fiber length, FC/APC connectors
Four of these items ordered yesterday from http://afwtechnologies.com.au/sm_coupler.html
Attached is the modifed Pentek whitening board schematic. It includes the yet to be installed 1nF capacitors and comments.
Here is the transfer function and cutoff frequency (pole) of the first stage low pass circuit of the Pentek whitening board.
R1 = R2 = 49.9 Ohm, R3 = 50 kOhm, C = 0.01uF. Given a differential voltage of 30 volts, the voltage across the 50k resistor should be 29.93 volts.
So low pass RC filter with one pole at 1 MHz.
I have updated the schematic, up to the changes mentioned by Rana plus some notes, see the DCC link here: [PLACEHOLDER]
I should have done this by hand...
There seems to be something funny going on with MATLAB's license authentication on the control room workstations. Earlier today, I was able to start MATLAB on pianosa, but now attempting to run /cvs/cds/caltech/apps/linux64/matlab/bin/matlab -desktop results in the message:
License checkout failed.
License Manager Error -15
MATLAB is unable to connect to the license server.
Check that the license manager has been started, and that the MATLAB client machine can communicate
with the license server.
Troubleshoot this issue by visiting:
License path: /home/controls/.matlab/R2013a_licenses:/cvs/cds/caltech/apps/linux64/matlab/licenses/license.dat:/cv
Licensing error: -15,570. System Error: 115
Thanks to some expertly timed coffee from Ignacio, I have been able to achieve indefnite locks of the DRMI, first on a 1F/3F mix (P:REFL11, S: REFL165, M:AS55), and then purely on 3F (P:REFL33, S:REFL165, S:REFL165). MICH is currently actuated on the ITMs.
I saved a snapshot of the current settings so I don't lose my settings. I think one thing that prevented earlier recipies from working is that whitening gains may have changed, which we don't typically note down when reporting input matrix settings
My current settings for 3F locking:
+30dB whitening gain, +136 demod phase
PRCL = 9 x I - 200 counts
+24dB whitening gain, +3 demod phase
SRCL = 1 x I, MICH = 5 x Q - 1000counts
MICH: G=-0.03; Acq FM4/5; Trig 2/3/6/9
PRCL: G=-0.003; Acq FM4/5; Trig 1/2/6/9
SRCL: G=0.2; Acq FM4/5; Trig 2/3/6/9
I've injected excitations into the control filter outputs via the LSC-FFC FMS (and notched the frequencies in the control filters themselves), and noted GPS times for offline sensing analysis. (Namely the 10 minutes following 1125398900)
Handing off to pure 3F was a little finicky at first, I needed to use some pretty large offsets in the MICH_B and PRCL_B FMs. (-1000 and -200 counts respectively). Once these offsets were found, the DRMI can acquire on 3F. Alignment is pretty important, too. Acquiring is much faster when the loop gains are "too high." i.e. I see a fair amount of gain peaking at ~300Hz. Nevertheless, things are stable enough as is that I didn't feel like digging into reducing the gains to quieter values.
Q and Ignacio were taking a second look at the Pentek interface board which we're using to acquire the POP QPD, ALS trans, and MCF/MCL channels. It has a differential intput, two jumper able whitening stages inside and some low pass filtering.
I noticed that each channel has a 1.5 kHz pole associate with each 150:15 whitening stage. It also has 2 2nd order Butterworth low pass at 800 Hz. Also there's a RF filter on the front end. We don't need all that low passing, so I started modifying the filters. Tonight I moved the 800 Hz poles to 8000 Hz. Tomorrow we'll move the others if Steve can find us enough (> 16) 1 nF SMD caps (1206 NPO).
After this those signals ought to have less phase lag and more signal above 1 kHz. Since the ADC is running at 64 kHz, we don't need any analog filtering below 8 kHz.
Since Andrey's SUS Drift mon screen back in 2007, we've had several versions which used different schemes and programming languages. Diego made an update back in January.
Today I added his stuff to the SVN since it was lost in the NFS disks somewhere. Its in SUS/DRIFT_MON/.
Since we've been updating our userapps directory recently to pull in the screens and scripts from the sites, we also got a copy of the Thomas Abbott drift mon stuff which is better (Diego actually removed the yellow/red functionality as part of the 'upgrade'), but more complicated. For now we have the old one. I updated the good values with all optics roughly aligned just a few minutes ago.
added the cron script for this to megatron to run at 8:44 AM each morning. Here's the new MegaCron attached :-()-
** it takes ~13 minutes to complete on megatron
# m h dom mon dow command
#0 */1 * * * bash /home/controls/public_html/summary/bin/c1_summary_page.sh > /dev/null 2>&1
#15 5 * * * /ligo/apps/nds2/nds2-megatron/test-restart
# MEDM Screen caps for the webpage
2,13,25,37,49 * * * * /cvs/cds/project/statScreen/bin/cronjob.sh
# op340m transplants -ericq
We investigated the ETMY oplev table set up and did not find a red herring.
Two 2 years vs one day plot below.
ps: thanks Q for fixing DTT, the auto scaling is not working at sampling rate 10 min and 1 hr period??
I experimented with removing somethings here and there to reduce the c1cal runtime. Eventually I deleted the LSC Sensing Matrix from it.
After removing sensing matrix, the run time is now down to 6 usec.
Back in 2011, JoeB wrote some entries on how to automatically update the Simulink webview stuff.
Somehow, the cron broke down over the years. I reran the matlab file by hand today and it worked fine, so now you can see the up to date models using the internet.
Today, I remeasured the transfer function for MC2 to MCL in order to improve the subtraction performance for MCL and to quantify just how precisely it needs to be.
Here is the fit, and the measured coherence. Data is also attached here: TF.zip
OMG, I forgot to post the data and any residuals. LOL!
The transfer function was fitted using vectfit with a weighting based on coherence being greater than 0.95.
I then used the following filters to do FF on MCL online:
Here are the results:
Performance has definelty increased when compared to previous filters. The reason why I think we still have poor performance at 3 Hz, is 1) When I remeasured the transfer function, Eric and I were expecting to see a difference on its shape due to the whitening filters that were loaded a couple days ago. 2) Assuming the transfer function is correct, there is poor coherence at 3 Hz 3) The predicted IIR subtraction is worst at this frequency.
ran the ON script several times and it kept pulling it away from good alignment, even when TRX was > 0.9.
Also, for what reason was this model run at 16 kHz?? Makes no sense to me to have a low frequency servo system run so high. Only makes for more digital precision noise, more CPU time, etc. Of course, running it at 2k would mean having to think about all of the AA filtering needed to go up/down from 2k to 16k.
The IMC often was making that scratchy noise when first catching lock and sometimes breaking. Thinking of the crappy crossover sit that EQ showed in his latest plots, I decided that it didn't make sense to acquire lock with an unstable PZT/EOM crossover, so I have changed mcdown to acquire with +13 dB Fast Gain and its much fast now and no longer makes that sound.
I also changed the caput command from 'caput -l' to 'caput -c -l' to see if the async 'wait for callback' feature will insure that the commands get sent. I witnessed the mcdown not actually writing all of its commands once or twice tonight. With the MC Boost left on its never going to lock.
mcdown has been committed to SVN. Please, if you have recently edit mcup and Autolock, commit them to the SVN or else I will delete them and do an svn up.
Let's order a pair of 35.5 MHz Wenzel for this guy and package like Rich has done for the WB low noise oscillators.
WE're only sending 6 dBm into it now and its using a 13 dBm mixer. Bad for PMC stability.
Also, if anyone has pix of the servo card, please add them to the DCC page for the PMC.
From the AFW website about our product, the POBC-64-C-1-7-2-25dB:
port1 slow axis -> port3 slow axis
port2 slow axis -> port3 fast axis
The promised historical comparisons follow. The crossover looks mostly the same as before. There is a new feature in the OLG at 50-60kHz; what could've changed about the EOM path in that time?
Maybe we just don't understand the splitter/combiners.
After an email from Eric G, I think this is the case.
If you read the text at Thorlabs about Fiber-Based Polarization Beam Combiners/Splitters, it suggests that these things take input beams both aligned to their slow axes, and outputs one field along the slow, and one orthogonal to it on the fast axis. Which is exactly what we don't want for a beat.
There has been some discussion here and there of using fiber coupled IR beats for ALS. A few weeks ago, and again today with Eric G, I poked around a bit with the fiber box Manasa set up for the FOL scheme.
Somehow, the IR beatnote is ~1000 times smaller than expected, both with the Thorlabs fiber coupled PD and a fiber coupled NF 1611.
In essence, after the fiber combiner, there is on the order of hundreds of uW each of PSL and AUX X IR light. The output of the fiber from each source looks nice and gaussian. The DC output of the 1611 indicates that it is seeing the right level of light. The green beatnote exists with good SNR at twice the IR beat frequency, so we know that the IR beat isn't some junky modes beating.
For the 1611, we would expect an RF signal of ~1mW*0.9A/W*700V/A -> .6V / +8dBm. Instead we see ~2mV / -40dBm.
Incidentally, there is some 20mV / -20dBm signal at ~400kHz, presumably from the green PDH modulation at ~200k.
(The level out of the thorlabs PD is similarly tiny; it doesn't have a DC output though, so we don't know the DC power that the active surface really sees. Not that I expect it to be much different, but the NF just makes it easier to estimate.)
The only things that should be able to cause the beat to be smaller than expected from the power levels are mode matching and polarization matching. All the fibers are single mode, so mode matching should be effectively 100%. Maybe somthing fishy is happening with the polarizations, but they'd have to be really maliciously close to orthogonal to cause this level of mismatch.
Maybe we just don't understand the splitter/combiners. Mysterious.
I took some transfer functions of the IMC loop and crossover, being careful that the PC drive never exceeding 1V during the measurements.
I then did some algebra to try and back out the individual loop paths, without having to make assumptions/approximations about the loop gain being high enough. This only really works in the region where both the open loop and crossover measurements have coherence.
It seems to me that the PZT path has pretty low phase margin on its own, but maybe this is ok, since its never really meant to run solo. The EOM path shape is harder to understand.
The data I took, and code that made the above plot is attached. This afternoon, I'll post an update comparing the measured OLG and crossover to earlier measurements.
To help find out if Steve really melted the inside of our precious seismometer, lets hook it up using the handheld seismo wand and see if it produces volts when we shake the ground.
Also, please stop using names like GurA or Gur1 or GurSuzy. We have GurX and GurY because they are at those ends. Anything else is confusing.
I moved Gur A from ETMX to ETMY . Gur B at ETMY was disconnected and its cable connected to Gur A
It seems that Gur A is alive. I will stop using A and B names after we stop swapping components.